# Remasking Discrete Diffusion Models with Inference-Time Scaling

### v2

@@ 2 Background
Discrete Diffusion Models  Diffusion models [47, 18] are characterized by parametric models $p_{\theta}$ trained to reverse a fixed noising process $q$ that maps clean data $\mathbf{x}$ to increasingly noisy latent variables $\mathbf{z}_{t}$, for $t\in[0,1]$. Austin et al. [2], Sahoo et al. [38], Shi et al. [45] extend this framework to discrete signals. Formally, we define $\mathbf{x}\in\{0,1\}^{|V|}\subset\Delta^{|V|}$ as one-hot vectors, where $|V|$ is the vocabulary size and $\Delta^{|V|}$ is the simplex over this vocabulary. For finite-time processes, we let $T$ be the number of time steps. For each time step $t(i)=\frac{i}{T}\in[0,1]$ (i.e., for $i=0,1,\ldots,T$), each intermediate latent variable $\mathbf{z}_{t(i)}\in\{0,1\}^{|V|}$ has the marginal:
@@
Masked Diffusion Language Models  A special case of this formulation, known as ‘masking’ or ‘absorbing state’ diffusion, sets the limiting distribution $\bm{\pi}=\bm{m},$ a one-hot representation of a special [MASK] token. With $s(i)=\frac{i-1}{T}$ as the time step directly preceding $t$, these posteriors take the form:
@@
Importantly, despite the recent success of MDLM [38] and other absorbing state discrete diffusion models, these models suffer from a key drawback, which we call the failure to remask property. Namely, the posterior formula in (2) implies that any unmasked token $\mathbf{z}_{t}\neq\bm{m}$ must remain unchanged throughout the entire denoising process, for all $s\leq t,$ which Sahoo et al. [38] formalize as $q(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x})=\mathrm{Cat}(\mathbf{z}_{s};\mathbf{z}_{t})$ if $\mathbf{z}_{t}\neq\bm{m}$. Once a token is decoded, this prediction is ‘locked-in’, a limitation that these diffusion models share with AR models.
As in Sahoo et al. [38], a denoising neural network $\mathbf{x}_{\theta}$ is used for parameterizing $p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})=q(\mathbf{z}_{s}|\mathbf{z}_{t},\mathbf{x}_{\theta}(\mathbf{z}_{t}))$ and trained using the following objective:
@@ 3 Remasking Diffusion Models
3.1 ReMDM Forward Processes
We define the complete forward process of ReMDM as:
@@ 3.2 Negative Evidence Lower Bound
$\mathcal{L}^{\sigma}=\mathbb{E}_{\mathbf{z}_{0}\sim q(\mathbf{z}_{0}\mid\mathbf{x})}\bigg{[}-\log p_{\theta}(\mathbf{x}\mid\mathbf{z}_{0})\bigg{]}+\mathbb{E}_{t\in\{\frac{1}{T},\ldots,1\}}\mathbb{E}_{\mathbf{z}_{t}\sim q(\mathbf{z}_{t}\mid\mathbf{x})}T\bigg{[}\frac{(1-\sigma_{t})\alpha_{t}-\alpha_{s}}{1-\alpha_{t}}\log(\mathbf{x}_{\theta}^{\top}\mathbf{x})\bigg{]}.$
Since $\alpha_{t}\leq\alpha_{s},$ each term in the second expectation increases monotonically in $\sigma_{t}$. When $\sigma_{t}=0$, for all $t$, we recover the MDLM objective from (3). This implies that the MDLM NELBO produces a tighter bound compared to ReMDM. Furthermore, since the diffusion loss term in (3.2) is simply a reweighted version of that in (3), one can presumably reuse weights from a model trained with (3), i.e., using different $\sigma_{t}$ at training and inference. Indeed, we find this to be a performant strategy in practice. Moreover, we find the performance (test perplexity) between models trained with the MDLM and the ReMDM objectives to be comparable (see Appendix E.1), further justifying our reuse of pretrained weights, with the added benefit of more flexible sampling unlocked by ReMDM.
@@ 4.1 Design Strategies for the Remasking Schedule $\sigma_{t}$
[…]
@@ 4.2 Design Strategies for ‘Turning On’ ReMDM
Switch  We choose some $t_{switch}\in(0,1]$ and in (9) we have $[t_{on},t_{off})=[t_{switch},0)$. We denote this strategy as “ReMDM-switch.”
@@ 4.3 Comparison with Discrete Predictor-Corrector Samplers
The ReMDM sampler is more general than DFM corrector, since only the former can accommodate noise schedules with constant $\alpha_{t}$ for some range $[t,t-\Delta t]$ where $\Delta t>0$.
@@ 5.1.1 Text Generation
Generative Perplexity vs. MAUVE It is common practice to report generative perplexity (Gen PPL.) as a quality metric for text generation. However, for corrector samplers such as ReMDM, we find this metric can be uninformative. In line with previous work [63], we observe that one can get arbitrarily low Gen PPL. by tuning corrector schedule hyperparameters (see Appendix E.2.2) to achieve low Gen PPL. while sacrificing sentence entropy/diversity, with an extreme case of generating a sequence of repetitive words like “love love love…" and getting a Gen PPL. of 1.5 (see Appendix F.1). Given this phenomenon, we choose MAUVE as the primary quality metric since MAUVE measures the balance between generative perplexity and diversity.
@@
As shown in Table 1, ReMDM scales favorably with inference-time compute, achieving a 15.62$\times$ MAUVE score compared to masked diffusion models and a 2.23$\times$ MAUVE score compared to MDLM with corrector samplers. In contrast, masked diffusion models scale poorly with $T$, and corrector sampler methods saturate when $T$ is large. For faster sampling where a more limited computational budget is available, ReMDM is able to maintain sample quality better than baselines.
Ablation: Function of Each Component  In Figure 3, we sequentially ablate each building block of ReMDM. Starting with MDLM, we see that each of our proposed sampling improvements increases the MAUVE score, with the largest improvement coming from the remasking ability of ReMDM.
Table 1: ReMDM improves sample quality in inference-time scaling and faster sampling on OWT. ReMDM outperforms state-of-the-art masked diffusion models (SEDD; Lou et al. [26], MDLM; Sahoo et al. [38]) and masked diffusion models with corrector samplers such as Forward-Backward (FB; Campbell et al. [5]) and Discrete Flow Matching (DFM; Gat et al. [13]) corrector samplers. indicates nucleus sampling. For each $T$, the best diffusion MAUVE score is bolded.
@@ 5.1.2 Image Generation
Experimental Setup  We test ReMDM’s class-conditioned image generation ability. Concretely, we use a pretrained MaskGiT [7] that was trained on ImageNet [8] samples with $256\times 256$ pixels. The images were patchified and flattened to a sequence of $L=256$ and encoded according to a codebook with $1024$ tokens [10]. See Chang et al. [7] for more details about the model and training settings. We compare ReMDM’s sampler to the original MaskGiT sampling and to MDLM. Note that MaskGiT does not follow an ancestral sampling method but rather relies on an effective heuristic where masked tokens are decoded proportional to model confidence. However, MaskGiT is still restricted by the failure to remask property. For each sampler, we conditionally generate 50,000 images, with class labels randomly chosen from the 1,000 ImageNet categories, and we measure sample quality using Fréchet Inception Distance (FID; Heusel et al. [16]) and Inception Score (IS; Salimans et al. [42]).
@@
Ablation: Max-capped vs. Rescaled schedules  In Appendix E.3, we present the full ReMDM parameter search and find that ReMDM-rescale slightly outperforms ReMDM-cap and that for both increasing their corresponding $\eta$ parameters leads to improved results.
@@ 5.2 ReMDM Improves Guidance
Experimental Setup  We follow the setup from Schiff et al. [43] to explore controlled small molecule generation. Specifically, we use the QM9 dataset [37, 36] of $\sim$133k molecules and their character-based SMILES string representations [55]. Sequences were tokenized using a regular expression method [44] and padded to a maximum length of $L=32.$ We use the D-CFG and D-CBG methods defined in Schiff et al. [43] to conditionally generate molecules with higher ring counts (greater ${}^{\text{th}}$ percentile in the original dataset). We vary the guidance strength $\gamma\in\{1,2,3,4,5\}$ and scale inference steps $T\in\{32,64,128\}$. Our baselines consist of an AR model guided using either CFG or the popular classifier-based FUDGE method [58], MDLM, and the uniform diffusion language model (UDLM) proposed in Schiff et al. [43], which was also introduced to alleviate the no remasking limitation of masked diffusion. For ReMDM sampling, we use pretrained MDLM models. We explore the various strategies defined in Section 4 (see Appendix E.4 for full search results). After generating 1,024 sequences, we use the RDKit library [21] to determine whether sequences are valid and compute novelty of the generated sequences (number of unique valid sequences that do not appear in the original QM9 dataset) and mean ring count of novel sequences.
Results  In Figure 4, we display the trade-off that comes from increasing guidance strength $\gamma$. We only visualize results for samples that had at least 50 novel sequences. For both forms of guidance, CFG and CBG, ReMDM outperforms AR and diffusion approaches, pushing the novelty-property maximization frontier beyond that of the baseline methods. Additionally, ReMDM scales favorably with more inference-time compute, seen by the curves for larger $T$ dominating those for smaller $T$. For ReMDM, we found the best strategy to be a combination of the ReMDM-rescale (with $\eta_{rescale}=0.9$) and ReMDM-conf schedules. For D-CFG, we use the loop strategy, with $t_{on}=0.25,t_{off}=0.125$. In addition to these findings, in Appendix E.4.1, we present experimental results for maximizing a different property of interest, drug-likeness (QED; Bickerton et al. [3]).
Figure 4: ReMDM improves steer-ability by extending the novelty-property maximization frontier. Controlled generation for ring count maximization on QM9 dataset, varying compute $T$ and guidance strength $\gamma$. (Left) Discrete classifier-free guidance (D-CFG). (Right) Discrete classifier-based guidance (D-CBG) and FUDGE for AR.
Ablation: Confidence-based Schedule  Ablations show that using ReMDM-rescale in conjunction with the confidence-based scheduler led to improved results (see Appendix E.4.3 for more details).
@@ 5.3 ReMDM Improves Benchmark Performance
Experimental Setup Recent diffusion large language models (dLLMs) such as LLaDA 8B [28] and Dream 7B [60] have demonstrated on par downstream task performance with autoregressive language models of the same parameter scale. However, these dLLMs still suffer from the failure to remask property. To test the effect of remasking sampling on dLLLMs, we take LLaDA 8B Instruct and apply DFM and ReMDM samplers (see Appendix D.4 for sampler details). Specifically, we use Countdown (bidirectional reasoning) [59] and TruthfulQA (factual knowledge grasp) [24]. In Countdown, the model is given four integers and asked to develop equations where first three integers produce the fourth. In TruthfulQA, the model is asked 817 questions on factual knowledge. We calculate the ROUGE [23] scores between the model and correct answers, as well as the model and plausibly wrong answers. The difference between the right and wrong ROUGE scores is reported. For both of the tasks, we use max-length $L=32$ and no semi-autoregressive sampling [28]. See Appendix D.4 & F.3 for more setting details.
Results  Since the metric gains are not as significant as in Table 1, we repeat experiments with multiple random seeds to report mean values and 95% confidence intervals. In Table 3, LLaDA with ReMDM consistently performs the best and exceeds the original LLaDA with statistical significance.
@@
LLaDA 45.2±${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\text{0.2}}}$ 27.1±${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\text{0.4}}}$ / 30.1±${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\text{0.4}}}$ / 27.2±${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\text{0.4}}}$
LLaDA-DFM 44.8±${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\text{0.2}}}$ 28.2±${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\text{0.4}}}$ / 31.1±${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\text{0.4}}}$ / 28.3±${}_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\pm\text{0.4}}}$

### v3

@@ 2 Background
Discrete Diffusion Models  Diffusion models [50, 18] are characterized by parametric models $p_{\theta}$ trained to reverse a fixed noising process $q$ that maps clean data $\mathbf{x}$ to increasingly noisy latent variables $\mathbf{z}_{t}$, for $t\in[0,1]$. Austin et al. [2], Sahoo et al. [41], Shi et al. [48] extend this framework to discrete signals. Formally, we define $\mathbf{x}\in\{0,1\}^{|V|}\subset\Delta^{|V|}$ as a one-hot vector representing a token, where $|V|$ is the vocabulary size and $\Delta^{|V|}$ is the simplex over this vocabulary. For finite-time processes, we let $T$ be the number of time steps. For each time step $t(i)=\frac{i}{T}\in[0,1]$ (i.e., for $i=0,1,\ldots,T$), each intermediate latent variable $\mathbf{z}_{t(i)}\in\{0,1\}^{|V|}$ has the marginal:
@@
Masked Diffusion Language Models  A special case of this formulation, known as ‘masked’ or ‘absorbing state’ diffusion, sets the limiting distribution $\bm{\pi}=\bm{m},$ a one-hot representation of a special [MASK] token. With $s(i)=\frac{i-1}{T}$ as the time step directly preceding $t$, these posteriors take the form:
@@
Following Sahoo et al. [41], we omit the subscript $(i)$ in $s(i)$ and $t(i)$ in the rest of this paper. Importantly, despite the recent success of MDLM [41] and other absorbing state discrete diffusion models, these models suffer from a key drawback, which we call the failure to remask property. Namely, the posterior formula in (2) implies that any unmasked token $\mathbf{z}_{t}\neq\bm{m}$ must remain unchanged throughout the entire denoising process. Formally, in masked diffusion models, equation 1 ensures that any token $\mathbf{x}$ only possesses two possible values: unmasked, i.e., keeping unchanged, and masked, i.e., $\mathbf{m}$, which means that in the reverse process, for any token $\mathbf{z}_{t}\neq\mathbf{m}$, for all $s\leq t$, plugging $\mathbf{z}_{t}=\mathbf{x}$ into equation 2 leads to $q(\mathbf{z}_{s}\mid\mathbf{z}_{t},\mathbf{x})=\mathrm{Cat}(\mathbf{z}_{s};\mathbf{z}_{t})$. Intuitively, once a token is decoded, this prediction is ‘locked-in’, a limitation that these diffusion models share with AR models.
As in Sahoo et al. [41], a denoising neural network $\mathbf{x}_{\theta}$ is used for parameterizing $p_{\theta}(\mathbf{z}_{s}|\mathbf{z}_{t})=q(\mathbf{z}_{s}|\mathbf{z}_{t},\mathbf{x}_{\theta}(\mathbf{z}_{t}))$ and trained using the following variational objective (see the derivation in Sahoo et al. [41]):
@@ 3 Remasking Diffusion Models
3.1 ReMDM Probability Decomposition
We define the probability decomposition of ReMDM as:
@@ 3.2 Negative Evidence Lower Bound
$\small\mathcal{L}^{\sigma}=\mathbb{E}_{\mathbf{z}_{0}\sim q(\mathbf{z}_{0}\mid\mathbf{x})}\bigg[-\log p_{\theta}(\mathbf{x}\mid\mathbf{z}_{0})\bigg]+\mathbb{E}_{t\in\{\frac{1}{T},\ldots,1\}}\mathbb{E}_{\mathbf{z}_{t}\sim q(\mathbf{z}_{t}\mid\mathbf{x})}T\bigg[\frac{(1-\sigma_{t})\alpha_{t}-\alpha_{s}}{1-\alpha_{t}}\log(\mathbf{x}_{\theta}^{\top}\mathbf{x})\bigg].$ (9)
Since $\alpha_{t}\leq\alpha_{s},$ each term in the second expectation increases monotonically in $\sigma_{t}$. When $\sigma_{t}=0$, for all $t$, we recover the MDLM objective from (3). This implies that the MDLM NELBO produces a tighter bound compared to ReMDM. Furthermore, since the diffusion loss term in (9) is simply a reweighted version of that in (3), one can presumably reuse weights from a model trained with (3), i.e., using different $\sigma_{t}$ at training and inference. Indeed, we find this to be a performant strategy in practice. Moreover, we find the performance (test perplexity) between models trained with the MDLM and the ReMDM objectives to be comparable (see Appendix E.1), further justifying our reuse of pretrained weights, with the added benefit of more flexible sampling unlocked by ReMDM.
@@ 4.1 Design Strategies for the Remasking Schedule $\sigma_{t}$
$\sigma_{t}^{(\ell)}=\eta_{conf}^{(\ell)}\cdot\sigma_{t}\text{~~, where~~}\eta_{conf}^{(\ell)}=\frac{\exp(-\psi_{t}^{(\ell)})}{\sum_{l=1}^{L}\exp(-\psi_{t}^{(\ell^{\prime})})}.$
@@ 4.2 Design Strategies for ‘Turning On’ ReMDM
Switch  We choose some $t_{switch}\in(0,1]$ and in (10) we have $t_{on},t_{off})=[t_{switch},0)$. We denote this strategy as “ReMDM-switch.”
@@ 4.3 Comparison with Discrete Predictor-Corrector Samplers
The ReMDM sampler is more general than DFM corrector, since only the former can accommodate noise schedules with constant $\alpha_{t}$ for some range $t,t-\Delta t]$ where $\Delta t>0$.
@@ 5.1.1 Text Generation
Generative Perplexity vs. MAUVE It is common practice to report generative perplexity (Gen PPL.) as a quality metric for text generation. However, for corrector samplers such as ReMDM, we find this metric can be uninformative. In line with previous work [68], we observe that one can get arbitrarily low Gen PPL. by tuning corrector schedule hyperparameters (see Appendix E.2.2) to achieve low Gen PPL. while sacrificing sentence entropy/diversity, with an extreme case of generating a sequence of repetitive words like “love love love…" and getting a Gen PPL. of 1.5 (see Appendix F.1). Given this phenomenon, we choose MAUVE as the primary quality metric since MAUVE measures the balance between generative perplexity and diversity (see Pillutla et al. [37] for a more detailed discussion).
@@
[…]
Ablation: Function of Each Component  In Figure 3, we sequentially ablate each building block of ReMDM and report the MAUVE score of samples. Starting with MDLM, we see that each of our proposed sampling improvements increases the MAUVE score, with the largest improvement coming from remasking ability of ReMDM.
Table 1: ReMDM improves sample quality in inference-time scaling and faster sampling on OWT. ReMDM outperforms state-of-the-art masked diffusion models (SEDD; [29], MDLM; [41]) and masked diffusion models with corrector samplers such as Forward-Backward (FB; [5]) and Discrete Flow Matching (DFM; [13]) corrector samplers. indicates nucleus sampling. For each $T$, the best diffusion MAUVE score is bolded.
@@ 5.1.2 Image Generation
Experimental Setup  We test ReMDM’s class-conditioned image generation ability. Concretely, we use a pretrained MaskGiT [7] that was trained on ImageNet [8] samples with $256\times 256$ pixels. The images were patchified and flattened to a sequence of $L=256$ and encoded according to a codebook with $1024$ tokens [10]. See Chang et al. [7] for more details about the model and training settings. We compare ReMDM’s sampler to the original MaskGiT sampling and to MDLM. Note that MaskGiT does not follow an ancestral sampling method but rather relies on an effective heuristic where masked tokens are decoded proportional to model confidence. However, MaskGiT is still restricted by the failure to remask property. For each sampler, we conditionally generate 50,000 images, with class labels randomly chosen from the 1,000 ImageNet categories, and we measure sample quality using Fréchet Inception Distance (FID; [16]) and Inception Score (IS; [45]).
@@
Ablation: Max-capped vs. Rescaled schedules  In Appendix E.3, we present the full ReMDM parameter search results. We find that ReMDM-rescale slightly outperforms ReMDM-cap, and that for both strategies, increasing their corresponding $\eta$ parameters leads to improved results.
@@ 5.2 ReMDM Improves Guidance
[…]
Results  In Figure 4, we display the trade-off that comes from increasing guidance strength $\gamma$. We only visualize results for samples that had at least 50 novel sequences. For both forms of guidance, CFG and CBG, ReMDM outperforms AR and diffusion approaches, pushing the novelty-property maximization frontier beyond that of the baseline methods. Additionally, ReMDM scales favorably with more inference-time compute, seen by the curves for larger $T$ dominating those for smaller $T$. For ReMDM, we found the best strategy to be a combination of the ReMDM-rescale (with $\eta_{rescale}=0.9$) and ReMDM-conf schedules. For D-CFG, we use the loop strategy, with $t_{on}=0.25,t_{off}=0.125$. In addition to these findings, in Appendix E.4.1, we present experimental results for maximizing a different property of interest, drug-likeness (QED; [3]).
Figure 4: ReMDM improves steer-ability by extending the novelty-property maximization frontier. Lines show controlled generation for ring count maximization on QM9 dataset, varying compute $T$ and guidance strength $\gamma$. Curves in the upper right region are better. (Left) Discrete classifier-free guidance (D-CFG). (Right) Discrete classifier-based guidance (D-CBG) and FUDGE for AR.
Ablation: Importance of using Confidence-based Schedule  Our parameter tuning revealed that using ReMDM-rescale in conjunction with the confidence-based scheduler led to improved results (see Appendix E.4.3).
@@ 5.3 ReMDM Improves Benchmark Performance
Experimental Setup Recent diffusion large language models (dLLMs) such as LLaDA 8B [31] and Dream 7B [64] have demonstrated on par downstream task performance with autoregressive language models of the same parameter scale. However, these dLLMs still suffer from the failure to remask property. To test the effect of remasking sampling on dLLLMs, we take LLaDA 8B Instruct and apply DFM and ReMDM samplers (see Appendix D.4 for sampler details). Specifically, we use Countdown (bidirectional reasoning) [63] and TruthfulQA (factual knowledge grasp) [27] as benchmarking tasks. In Countdown, the model is given four integers and asked to develop equations where first three integers produce the fourth using addition, subtraction, multiplication and division operations. In TruthfulQA, the model is asked 817 questions on factual knowledge. We calculate the ROUGE [26] scores between the model and correct answers, as well as the model and plausibly wrong answers. The difference between the right and wrong ROUGE scores is reported. For both of the tasks, we use max-length $L=32$ and no semi-autoregressive sampling [31]. See Appendix D.4 & F.3 for more setting details.
Results  We repeat experiments with multiple random seeds to report mean values and 95% confidence intervals. In Table 3, LLaDA with ReMDM consistently performs the best and exceeds the original LLaDA with statistical significance.
@@
[…]